OpenStudy (danielbarriosr1):

Adults tickets to a play cost $22. Tickets for children cost $15. Tickets for a group of 11 people cost a total of $228. Write and solve a system of equations to find how many children and how many adults were in the group.

OpenStudy (danielbarriosr1):

Please help

OpenStudy (anonymous):

Adult tickets x numbers child tickets y numbers 22 x + 15 y = 228 .............1 1 x + 1 y = 11 .............2 Eliminate y multiply (1)by 1 Multiply (2) by -15 22 x 15 y = 228 -15 x -15 y = -165 Add the two equations 7x =63 /7 x=9.00 OR 9 plug value of x in (1) 22 x + 15 y = 228 198 + 15 y = 228 15 y = 228-198 15 y = 30 y = 2.00 OR Adult tickets 9 numbers child tickets 2 numbers

OpenStudy (danielbarriosr1):

Thanks again @sunnyshores

OpenStudy (anonymous):

np!